OpenStudy (anonymous):

6 divided by (5 - √3)

OpenStudy (anonymous):

I got 30 + 6√3 divided by 22, but that's not one of my choices

OpenStudy (anonymous):

yay geerky is here

geerky42 (geerky42):

Multiply both numerator and denominator by the conjugate of the denominator.

OpenStudy (anonymous):

i did that originally and got 30 + 6√3 divided by 25 - √9

geerky42 (geerky42):

Try to simplify it.

OpenStudy (anonymous):

√9 simplified is 3 so i can simplify the expression to 30 + 6√3 divided by 22

geerky42 (geerky42):

OK, now simplify 30 + 6√3 divided by 22

OpenStudy (anonymous):

i dont know how to simplify that

geerky42 (geerky42):

\[\frac{30 + 6\sqrt{3}}{22} = \frac{30 + 6\sqrt{3}}{2 \cdot 11} = \frac{30}{2 \cdot 11} + \frac{6\sqrt{3}}{2\cdot 11} = \frac{15}{11} + \frac{3\sqrt{3}}{11} = \boxed{\frac{15 + 3\sqrt{3}}{11}}\]

OpenStudy (anonymous):

is that 3√3 in the box?

geerky42 (geerky42):

\[\Huge \boxed{\frac{15 + 3\sqrt{3}}{11}}\]

geerky42 (geerky42):

Yes.

geerky42 (geerky42):

Is this clear?

OpenStudy (anonymous):

oh so you have to take 2 and 11 out of the expression since its the GCF?

OpenStudy (anonymous):

yeah i get it now, thanks a lot!

geerky42 (geerky42):

Take 2 out, yeah.

geerky42 (geerky42):

No problem, glad I helped.